OpenStudy (anonymous):

HELP PLEASE!!! Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution x - y = 3 and x + y = 7 Solve the system by using elimination or elimination with multiplication. One point for the steps and one point for the correct solution. 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (cornzyblack):

OK i can just make x the subject of formula in both equations equation 1----- x = 3 + y equation 2----- x = 7 - y then i equate them so that 3 + y = 7 - y collect like terms 2y = 7 - 3 y = 4/2 = 2 Substituting in eqn 1 we have x = 3 + 2 = 5 x = 5, y = 2

OpenStudy (kinggeorge):

Well for the first equation, simply add the two equations together. After that, you get \(2x = 10\) so \(x=5\) and then solving for \(y, \quad y=2\). As for the second system, multiply the first by 3 and the second by 4 to get\[\begin{matrix} 9x+12y=-6 \\ 8x-12y=40 \end{matrix} \]and add them together again to get \(17x=34\) so here \(x=2\). Solving for \(y, \quad y=-2\)